media%2F19b%2F19b2dcb8-2301-4459-b8d5-0c
Answer #1

(A) d = u t + a t^2 / 2

0.811 = 0 + a (2.52^2 / 2)

a = 0.255 m/s^2


(b) m2g - T2 = m2 a

0.500 x 9.81 - T2 = 0.500 x 0.255

T2 = 4.777 N

(c) T1 - m1g = m1 a

T1 = 0.460 ( 9.81 + 0.255) = 4.62 N

(d) alpha = a / R = 0.255 / 0.05 = 5.10 rad/s^2


(e) net torque = R( T2 - T1) = 7.85 x 10^-3 N m

net torque = I alpha

7.85 x 10^-3 = I ( 5.10)

I = 1.54 x 10^-3 kg m^2